# In Search of Lost Online Test-time Adaptation: A Survey

### v2

@@ 4.2 Is OTTA still working with ViT?
We evaluate the CIFAR-10-C and CIFAR-10.1 datasets with batch sizes $1$ and $16$ and show the result in Fig. 8. To clearly understand the pattern of predictions, we discuss our observations in three aspects: 1) variation in corruption types, 2) variation in batch sizes, and 3) variation in adaptation strategies.
Corruption Type. As depicted in Fig. 8 and Fig. 15, a notable observation is that most methods experience high error rates with noise corruptions, irrespective of their batch size. However, these methods exhibit reasonable performance for some other types of corruption. This difference may be attributed to the random and unpredictable nature of noise corruptions, as opposed to more structured types like snow, zoom, or brightness, which are potentially more comprehensible for online adaptation.
Furthermore, adapting to noise corruption poses a significant challenge for confidence optimization-based methods (Sec. 3.1.3), regardless of the batch size. This difficulty can be linked to the substantial domain gap and the unpredictable nature of noise patterns discussed earlier. Although these strategies aim to increase the model’s confidence, they are not equipped to directly correct erroneous predictions.
Batch Size. Variations in batch size do not significantly alter the mean error, with notable exceptions for Tent, Conj-CE, and MEMO. As observed in Sec. 4.4, we conclude that for purely optimization-based methods, a larger batch size can stabilize loss optimization, thereby benefiting adaptation. However, incorporating consideration of prediction reliability can substantially mitigate the constraints imposed by smaller batch sizes, as evidenced in methods like CoTTA, TAST, and SAR. A similar pattern is observed in CIFAR-10.1, where two entropy-based methods (Tent and Conj-CE) exhibit limitations at small batch sizes.
Adaptation Strategy. SAR and RoTTA demonstrate stable performance regardless of domain or batch size variations. The memory bank in RoTTA contributes to maintaining global information, rendering it more batch-agnostic. From a different angle, SAR achieves flat minima, which ensures model optimization for stability and prevents biased learning during adaptation. MEMO also displays impressive performance in certain domains, even with batch sizes as small as $1$.
4.2.2 On CIFAR-100-C Benchmark
Figure 9: Comparison of the OTTA performance on the CIFAR-100-C severity level $5$. The upper and bottom plots show the experiments conducted with batch size $16$ and $1$ based on the LayerNorm updating strategy, respectively.
The performance on the CIFAR-100-C dataset exhibits a trend similar to that observed in the CIFAR-10-C dataset. To ensure a concise and focused discussion, we will explore specific adaptation strategies only when their performance patterns distinctly differ from those of the CIFAR-100-C dataset.
A notable observation is the comparatively less robust performance on CIFAR-100-C, especially when using a batch size of $16$. This decrease in performance is likely due to the greater complexity and variety in the CIFAR-100-C dataset, which includes a more extensive range of classes.
Adaptation Strategy. CoTTA demonstrates a noticeable decrease in effectiveness across most types of corruption, particularly with a batch size of $16$. This performance reduction can be partly attributed to the significant number of classes, for instance, comparing $10$ in CIFAR-10-C to $100$ in CIFAR-100-C. Additionally, the stochastic parameter reset might lead to the loss of newly acquired knowledge about these varied domains. Similarly, SAR begins to exhibit limitations, especially in the presence of noise distortions such as Gaussian, shot, and impulse noise.
An interesting observation is the performance degradation in the contrast corruption domain. Here, RoTTA stands out as the only method that consistently outperforms direct inference, irrespective of batch size. This underscores the importance of effectively preserving valuable sample information, particularly for tackling batch-sensitive and challenging adaptation tasks. A similar trend is observable in Fig. 10.
4.2.3 On Imagenet-C Benchmark
Figure 10: Comparison of the OTTA performance on the ImageNet-C severity level $5$. The upper and bottom plots show the experiments conducted with batch size $16$ and $1$ based on the LayerNorm updating strategy, respectively
Adaptation Strategy. For the ImageNet-C dataset depicted in Fig. 10, when the batch size is set to 16, SAR, Conj-CE, and RoTTA outperform the source-only model in terms of mean error. In contrast, Tent, MEMO, and CoTTA demonstrate significantly poor results. The error rate within each domain exhibits a similar trend. Notably, Conj-CE, which conducts an additional inference for each batch for final prediction, markedly surpasses Tent across most domains and in mean error. This suggests a significant inter-batch distribution shift in ImageNet-C, such as style or class differences, indicating that strategies based on optimizing the current batch for predicting the next batch are less effective. Additionally, MEMO faces challenges in scenarios characterized by complex label sets and significant data diversity. The parameter reset in CoTTA might also adversely affect the model’s discriminative capability, especially in complex environments.
When the batch size is reduced to 1, only RoTTA maintains its performance level, implying that typical unsupervised loss functions may be inadequate for complex adaptation tasks. Concurrently, preserving valuable data can significantly reduce the performance disparity caused by domain shifts.
Batch Size. Compared to CIFAR-10-C and CIFAR-100-C, ImageNet-C experiences more pronounced performance degradation, particularly when the batch size is reduced to 1. This observation, combined with our analysis of adaptation strategies, suggests that datasets with higher complexity and difficulty are more sensitive to changes in batch size.
Corruption Type. Relative to CIFAR-10-C and CIFAR-100-C datasets (as shown in Fig. 8 and Fig. 9), a distinctive feature of the ImageNet-C dataset is the narrower performance gap across various types of corruption. This implies that datasets with fewer classes are likely to show greater performance variability in response to different corruptions. For ImageNet-C, encompassing 1,000 classes, we observe consistently poor performance across all corruption domains, likely due to substantial inter-batch class differences, which hinder effective learning.
4.2.4 On CIFAR-10-Warehouse Benchmark
We assess OTTA techniques on the newly introduced CIFAR-10-Warehouse dataset, which shares the same label set as CIFAR-10-C. For our evaluation, we selected two representative domains within CIFAR-10-Warehouse. These domains were specifically chosen to gauge the performance of OTTA methods under two distinct distribution shifts: the real-world shift and the diffusion synthesis shift. The Google split consists of images sourced from the Google search engine. This subset serves as a critical benchmark for evaluating the capability of contemporary OTTA methods in managing real-world distribution shifts. We assess OTTA performance across its 12 subdomains labeled G-01 to G-12. Each subdomain represents images predominated by different main colors, providing a diverse range of visual scenarios to test the adaptability and effectiveness of OTTA methods under real-world conditions.
Figure 11: Comparison of the OTTA performance on the Google split of CIFAR-10-Warehouse. The upper and bottom plots show the experiments conducted with batch size $16$ and $1$ based on the LayerNorm updating strategy.
Batch Size. Regarding the batch size differences depicted in Fig. 11, we observe that when the batch size is $16$, most OTTA methods match or surpass the performance of direct inference. This outcome suggests that the current OTTA methods are generally effective. Additionally, the performance of most OTTA methods remains stable when the batch size is reduced to $1$. However, methods such as Tent and Conj-CE exhibit performance degradation in most domains. This may be attributed to the instability of optimization for single-sample batches, especially in Tent, which focuses solely on optimizing entropy.
Adaptation Strategy. RoTTA and SAR demonstrate exceptional stability regardless of batch sizes. This stability is achieved by preserving high-quality data information in RoTTA and seeking flat minima in its optimization in SAR. We compare Conj-CE with Conj-Poly, where Conj-Poly refers to the adaptation strategy when the source pre-training loss is poly loss (Leng et al, 2022). In our experiment, we modified the adaptation strategy without altering the source pre-training loss to observe performance differences. Interestingly, even when the batch size was set to 1, and the source pre-trained loss was cross-entropy loss (where Conj-Poly would not be the presumed optimal choice), Conj-Poly still managed to outperform Conj-CE in terms of mean error. This finding challenges the conclusions drawn in the original Conj-PL paper, suggesting that Conj-Poly might be more effective than initially thought, even when not aligned with the source pre-training loss.
Diffusion split. We evaluate OTTA methods within the diffusion subdomain.
Figure 12: Comparison of the OTTA performance on the diffusion split of CIFAR-10-Warehouse. All the experiments are conducted with batch size $16$ based on the LayerNorm updating strategy. The dotted line indicates the baseline performance without any adaptation.
(a) Batch Size=16
(b) Batch Size=1
Figure 13: Mean error vs Giga Floating Point Operations Per Second (GFLOPs). All experiments are conducted on a severity level of $5$. We plot their correlation under batch size $16$ and $1$ settings for the CIFAR-10-C dataset.
Subdomain Type. The results presented in Fig. 12 exhibit a consistent trend across 12 subdomains, with the exception of DM-$05$. In DM-$05$, confidence optimization methods perform less effectively compared to direct inference. This can be attributed to the substantial domain shift leading to distorted data interpretation. Relying solely on optimization in these methods may exacerbate model biases, especially due to prediction inaccuracy. In contrast, these methods excel in other domains, indicating the varying domain gaps across each diffusion-based subdomain.
Adaptation Strategy. Another noteworthy observation is the stable performance of CoTTA, SAR, and RoTTA. By employing sharpness-aware minimization, SAR enables the model to reach a region in the optimization landscape that is less sensitive to data variations, resulting in stable predictions. CoTTA’s parameter reset strategy effectively mitigates biased adaptation, allowing for partial knowledge restoration from the source domain, contributing to its consistent performance, even in the challenging DM-$05$ subdomain. Lastly, RoTTA, utilizing an informative memory bank, achieves commendable performance across the subdomains.
Conclusion. From our extensive experiments, most OTTA methods display similar behavioral patterns across various datasets. This consistency underscores the potential of contemporary OTTA techniques in effectively managing diverse domain shifts. Particularly noteworthy are RoTTA and SAR, highlighting the significance of optimization insensitivity and information preservation.
4.3 Is OTTA efficient?
To evaluate the performance of OTTA algorithms, particularly under hardware constraints, we utilize GFLOPs as a metric, as illustrated in Fig. 13. Lower GFLOPs and mean error is preferable. Our observation shows that MEMO achieves high performance but incurs higher computational costs. In contrast, RoTTA successfully balances low error rates with efficient updates. This also suggests that reducing the batch size could be beneficial in achieving a balance between performance and computational efficiency.
Figure 14: Impact of varying batch sizes on CIFAR-10.1, CIFAR-10-Warehouse, CIFAR-100-C, ImageNet-C. The base model is Tent optimized on LayerNorm.
4.4 Is OTTA sensitive to Hyperparameter Selection?
Batch Size Matters, but only to an extent. Fig. 15 examines the impact of different batch sizes on Tent’s performance in the CIFAR-10-C dataset. It reveals that performance significantly varies with batch sizes from $1$ to $16$ across most corruptions. However, this variability diminishes with larger batch sizes ($16$ to $128$), indicating a reduced influence of LayerNorm updating on batch size, in contrast to traditional BatchNorm settings. This pattern is consistent across other datasets, as depicted in Fig. 14. Nevertheless, batch size remains crucial for stabilizing the optimization process. For example, a batch size of $16$ outperforms a batch size of $1$ in confidence optimization methods in the Google split of the CIFAR-10-Warehouse dataset. However, Larger batch sizes are essential for complex datasets like CIFAR-100-C and ImageNet-C, where direct inference struggles, emphasizing the need to tailor batch sizes to the data complexity. Additionally, Fig. 15 suggests that increasing batch sizes is not effective against challenging corruptions, such as Gaussian and Shot noise. This highlights the necessity for more advanced adaptation strategies beyond mere batch size adjustments in complex learning scenarios.
Table 2: The Comparison of model update strategies across five benchmarks: Layernorm only v.s. full model. $*{}^{*}$ indicates the variant of CoTTA with no parameter reset mechanism.
CoTTA*** 0.2468 0.7708 -67.98% 0.0750 0.1640 -54.27% 0.4682 0.8961 -47.75% 0.7229 0.9054 -20.16% 0.1820 0.5463 -66.68%
Optimization Layer Matters! To assess the critical role of LayerNorm, we compare LayerNorm-based optimization with full model optimization, as summarized in Table. 2. This ablation study primarily focused on CoTTA and MEMO, evaluating the impact of optimizing LayerNorm alone. A notable observation is, for all methods, LayerNorm update plays an important role in gaining high performance, underscoring its effectiveness in boosting model performance by avoiding significant forgetting of the source knowledge.
5 Future Directions
  • Refining OTTA in Realistic Settings: Future OTTA methods should undergo testing in realistic environments, employing advanced architectures, practical testbeds, and reasonable batch sizes. This approach aims to gain deeper and more relevant insights.
6 Conclusion
In this survey, we thoroughly examine Online Test-Time Adaptation (OTTA), detailing existing methods, relevant datasets, benchmarks for evaluation, and their implementations. Comprehensive experiments are conducted to evaluate the effectiveness and efficiency of current OTTA methods when applied to vision transformers. Our observations indicate that noise-synthesized domain shifts often present more significant challenges compared to other types of shifts, such as those encountered in real-world scenarios or diffusion environments. Additionally, the presence of a large number of classes in a dataset can lead to noticeable discrepancies between batches, potentially impacting OTTA models’ ability to retain consistent knowledge. This may result in learning difficulties and an increased risk of severe forgetting. To address these challenges, we find that updating the normalization layer with a memory bank or optimization flatness, combined with proper batch size selection, can effectively stabilize the adaptation process and mitigate forgetting. We hope this survey will serve as a foundational reference, providing valuable insights for researchers and practitioners keen to explore the dynamic field of OTTA.
Data Availability Statement: The data that support the findings of this study are openly available as summarized in Table 1. This open-soured repository [^3] https://github.com/Jo-wang/OTTA_ViT_survey includes all the implementation code and experimental configurations of this work.
Appendix A Appendix
This appendix offers tables about the categorization of OTTA methods and experiments conducted on the seven selected methods across the chosen datasets.
  • Table 3: The summary of existing OTTA methods.
  • Table 4: Experiments on the CIFAR-10 $\rightarrow$ CIFAR-10-C task.
  • Table 5: Experiments on the CIFAR-100 $\rightarrow$ CIFAR-100-C task.
  • Table 7: Experiments on the CIFAR-10 $\rightarrow$ CIFAR-10-Warehouse task.
[…]
Method Glass Fog Defoc. Impul. Contr. Gauss. Elastic. Zoom Pixel Frost Snow JPEG Motion Shot Brit. Mean GFLOPs
Source Only 34.60 14.14 12.20 39.47 38.95 60.12 14.70 10.61 25.26 10.39 7.67 16.51 15.30 55.17 3.26 23.89 269.81
Batch size = 16
Tent 49.02 6.81 5.29 65.93 11.91 84.93 10.96 5.12 10.41 6.68 5.96 14.21 8.25 85.60 2.43 24.90 269.81
CoTTA-LN 34.50 14.57 12.97 43.96 50.01 58.04 15.05 11.09 22.92 10.43 7.75 16.46 15.92 53.65 3.27 24.71 809.42
CoTTA-ALL 71.68 51.28 38.25 82.60 79.30 85.57 31.22 16.51 39.95 36.75 37.38 37.37 81.00 88.54 9.72 52.47 809.42
CoTTA***-LN 34.58 14.60 13.02 43.71 49.71 58.00 15.03 11.09 22.96 10.45 7.67 16.53 15.93 53.68 3.24 24.68 809.42
CoTTA***-ALL 77.17 78.07 70.48 85.98 81.41 86.57 74.54 66.80 68.95 78.34 72.30 79.23 79.62 88.83 67.91 77.08 809.42
SAR 34.39 14.13 12.18 39.38 34.92 60.67 14.70 10.61 25.17 10.39 7.67 16.51 15.28 55.59 3.26 23.66 539.61
Conj-CE 61.26 6.36 4.89 80.12 7.94 87.96 10.92 4.20 8.43 5.32 5.33 12.71 7.18 86.63 2.27 26.10 539.61
MEMO-LN 26.15 9.08 6.17 65.80 9.58 77.68 12.48 5.29 9.14 6.59 5.71 14.45 7.71 59.15 2.49 21.16 809.42
RoTTA-LN 34.47 13.72 11.70 38.72 37.35 59.55 14.29 10.14 25.38 10.25 7.42 16.31 14.70 54.59 3.07 23.44 269.81
TAST 81.25 88.58 81.23 83.31 87.40 84.15 78.54 79.59 78.03 81.12 79.51 78.92 81.93 83.55 75.00 81.47 268.45
Batch size = 1
Tent 70.90 55.50 39.38 87.28 78.20 77.92 25.22 30.77 26.38 71.19 11.85 57.65 69.19 76.82 5.48 52.25 16.86
CoTTA-LN 34.09 16.63 14.79 43.05 45.14 57.63 15.39 13.23 21.11 10.15 8.25 16.88 17.68 52.36 4.07 24.70 50.59
CoTTA-ALL 87.81 89.90 86.13 89.17 88.82 89.55 84.91 87.77 86.30 87.42 89.17 88.67 89.06 89.22 81.65 87.70 50.59
CoTTA***-LN 35.88 18.35 16.96 46.27 47.31 58.21 17.49 15.12 20.95 10.78 9.60 18.15 19.72 52.15 4.83 26.12 50.59
CoTTA***-ALL 89.79 89.76 86.30 89.63 89.66 89.78 89.17 89.12 89.32 88.04 89.44 89.18 88.76 89.67 87.83 89.03 50.59
SAR 34.63 14.04 12.03 39.55 37.62 61.17 14.63 10.61 24.85 10.38 7.62 16.63 15.20 55.67 3.27 23.86 33.73
Conj-CE 48.90 55.11 14.09 88.28 84.49 89.11 14.39 29.01 20.95 13.70 27.84 38.25 12.71 88.95 5.43 42.08 33.73
MEMO-LN 85.91 5.55 4.77 87.84 3.76 89.33 11.54 3.70 5.52 4.08 4.36 12.28 6.01 89.32 1.65 27.71 556.48
RoTTA-LN 34.46 13.72 11.71 38.72 37.36 59.55 14.29 10.12 25.39 10.24 7.43 16.32 14.71 54.62 3.07 23.45 16.86
TAST 81.23 88.59 81.35 83.59 87.76 84.37 78.78 79.32 77.96 81.20 79.61 78.98 82.05 83.63 75.48 81.59 16.78
@@ Appendix A Appendix
CoTTA-ALL 96.95 93.78 95.34 98.9 97.82 98.86 92.47 93.33 96.36 95.66 98.42 95.79 96.81 98.68 98.04 96.48
CoTTA***-LN 66.87 46.04 42.94 85.11 71.35 82.93 39.81 35.38 38.61 30.09 27.39 45.03 44.50 79.71 11.59 49.82
CoTTA***-ALL 98.62 98.36 98.26 98.87 98.6 98.95 98.11 97.64 98.63 98.25 98.62 98.38 98.04 98.90 98.49 98.45
@@
RoTTA-LN 62.22 33.34 30.92 76.11 58.78 75.41 28.51 24.40 39.63 24.10 19.55 35.49 31.59 72.36 7.14 41.30
TAST 95.77 97.82 95.97 98.10 98.11 97.96 94.85 94.93 94.04 95.99 95.13 96.33 95.63 97.73 93.25 96.11
@@
Batch size = 16
CoTTA-LN 77.06 68.50 75.4 92.10 87.60 92.06 63.28 67.24 50.28 71.36 80.74 52.46 71.58 89.46 46.40 72.37
CoTTA***-ALL 90.30 95.84 95.08 99.28 95.24 99.12 87.9 88.2 90.20 94.84 97.50 68.04 95.82 98.68 62.08 90.54
CoTTA***-LN 77.26 68.6 75.32 91.58 87.08 92.28 63.3 67.18 50.72 70.74 80.64 52.02 71.48 89.72 46.46 72.29
Conj-CE 68.30 42.96 76.74 93.70 84.52 96.60 34.46 45.64 31.00 83.90 55.12 35.28 53.32 84.88 28.6 61.00
Conj-Poly 68.26 47.48 54.08 93.68 75.82 96.68 34.08 47.42 29.94 81.26 65.06 36.06 68.36 87.7 29.16 61.00
RoTTA-LN 73.50 63.24 65.14 78.40 89.58 79.02 53.10 61.94 39.12 60.34 63.24 44.62 60.80 76.88 33.12 62.80
TAST 99.76 99.64 99.74 99.68 99.90 99.64 99.46 99.30 99.12 99.46 99.32 98.78 99.58 99.36 98.44 99.41
Batch size = 1
CoTTA***-ALL 99.62 99.70 99.54 99.60 99.82 99.80 98.82 99.48 99.30 99.28 99.58 99.20 99.60 99.64 98.70 99.45
CoTTA***-LN 82.84 72.14 84.56 94.98 92.66 94.66 66.16 70.64 48.88 78.10 83.78 51.62 76.08 91.92 45.12 75.61
RoTTA-LN 73.46 63.26 65.10 78.40 89.58 79.04 53.14 61.94 39.12 60.36 63.18 44.64 60.86 76.88 33.12 62.81
TAST 99.76 99.68 99.70 99.66 99.88 99.64 99.42 99.26 99.18 99.46 99.32 98.74 99.60 99.38 98.42 99.41
@@
Google G-01 G-02 G-03 G-04 G-05 G-06 G-07 G-08 G-09 G-10 G-11 G-12 Mean
Source-Only 20.41 12.42 13.39 11.53 19.80 17.74 26.51 26.4 23.12 17.36 15.61 11.37 17.97
Batch size = 16
Tent 19.23 11.08 13.59 11.07 18.7 17.07 27.25 26.63 25.17 16.8 15.35 9.80 17.65
CoTTA***-LN 20.55 12.48 13.54 11.47 20.04 17.93 27.25 26.71 23.30 17.81 15.88 11.41 18.20
CoTTA***-ALL 40.13 60.93 61.92 33.15 50.54 48.50 56.21 58.46 61.05 61.48 59.86 63.37 54.63
RoTTA-LN 20.29 12.37 13.39 11.51 19.73 17.58 26.40 26.40 23.09 17.30 15.52 11.35 17.91
TAST 66.74 71.80 62.10 76.22 79.94 80.28 84.20 76.02 77.67 83.63 83.02 76.23 76.49
Batch size = 1
CoTTA***-LN 20.70 13.50 14.11 11.97 21.35 19.31 27.29 28.18 24.57 18.49 15.73 12.33 18.96
CoTTA***-ALL 86.81 85.35 86.21 87.08 88.44 87.41 72.75 90.88 84.03 89.93 85.44 85.36 85.81
RoTTA-LN 20.29 12.37 13.36 11.51 19.73 17.56 26.44 26.40 23.09 17.30 15.52 11.37 17.91
TAST 66.86 72.10 62.32 76.33 79.99 80.37 84.28 75.51 77.77 83.75 83.21 76.50 76.58

### v3

@@ 4.2 Does OTTA still work with ViT?
We evaluate the CIFAR-10-C and CIFAR-10.1 datasets with batch sizes $1$ and $16$ and show results in Fig. 8. We discuss our observations in three aspects.
Corruption Types. Across both batch sizes, most methods exhibit higher error rates in response to noise-induced corruptions, as illustrated in Fig. 8 and Fig. 13. In contrast, these methods tend to perform better when dealing with structured corruptions, such as snow, zoom, or brightness. This pattern suggests that various corruption types reflect different degrees of divergence from the corruption domain to the source dataset. Especially, adapting to noise corruption poses a significant challenge for confidence optimization-based methods (such as Tent, MEMO-LN, and Conj-CE), regardless of the batch size. This difficulty can be linked to the substantial domain gap and the unpredictable nature of noise patterns discussed earlier. Although these strategies aim to increase the model’s confidence, they are not equipped to directly correct erroneous predictions. It is worth noting that ROID surpasses all other baselines on noise-based corruptions. This may attributed to the certainty- and diversity-weighted SCE loss, which is designed to address noise issues while avoiding trivial solutions. CIFAR 10.1 stands out as an exception, as it is not associated with any specific corruption. Instead, it represents real-world data that share the same label set as CIFAR 10. Compared to other corruption domains — except the brightness domain — CIFAR 10.1 exhibits a smaller domain gap, as evidenced by the lower error rate achieved through direct inference. In this context, most methods yield results comparable to direct inference. Notably, Conj-CE excels with a batch size of 16, while CoTTA*-LN, CoTTA-LN, and SAR stand out when the batch size is reduced to 1.
Batch Sizes. Batch size is found to aid pure optimization-based methods like Tent, Conj-CE, and MEMO-LN in outperforming direct inference (i.e., source only). Similarly, batch size significantly influences the performance of ROID. It achieves the highest performance, with a significant margin over others, in terms of mean error on CIFAR-10-C when the batch size is 16. Nevertheless, it makes almost entirely erroneous predictions under a batch size of 1. As detailed in Section 4.4, our analysis leads to the conclusion that larger batch sizes tend to stabilize loss optimization, which is beneficial for adaptation processes in these methods. This phenomenon is also evident in CIFAR-10.1, where entropy-based methods such as Tent and Conj-CE face challenges at smaller batch sizes.
However, methods that incorporate prediction reliability can effectively navigate the limitations associated with smaller batch sizes. For instance, SAR capitalizes on sharpness-aware minimization to enhance its robustness. Similarly, CoTTA-LN employs selective augmentation in tandem with a reset strategy, averting erroneous optimization paths. In addition, RoTTA employs a well-designed memory bank to ensure batch-level stability. Conversely, TAST underperforms compared to the source-only baseline, even when utilizing larger batch sizes. This could be due to its methodology being specifically tailored for the ResNet family, which may not translate as effectively to other architectures.
@@
CoTTA-ALL 96.95 93.78 95.34 98.90 97.82 98.86 92.47 93.33 96.36 95.66 98.42 95.79 96.81 98.68 98.04 96.48
CoTTA*-LN 66.87 46.04 42.94 85.11 71.35 82.93 39.81 35.38 38.61 30.09 27.39 45.03 44.50 79.71 11.59 49.82
CoTTA*-ALL 98.62 98.36 98.26 98.87 98.60 98.95 98.11 97.64 98.63 98.25 98.62 98.38 98.04 98.90 98.49 98.45
@@
RoTTA 62.22 33.34 30.92 76.11 58.78 75.41 28.51 24.40 39.63 24.10 19.55 35.49 31.59 72.36 7.14 41.30
TAST 70.52 65.33 58.90 79.33 84.50 76.18 51.21 52.46 49.93 46.23 41.92 49.08 59.31 72.95 20.79 58.58
ROID 98.99 98.99 98.99 99.00 98.99 99.00 98.99 98.99 98.99 99.00 99.00 98.99 98.99 99.00 99.00 98.99
4.2.2 On CIFAR-100-C Benchmark
The performance on the CIFAR-100-C dataset exhibits a trend similar to that observed in the CIFAR-10-C dataset, as shown in Table 2.
@@ 4.2.2 On CIFAR-100-C Benchmark
4.2.3 On Imagenet-C Benchmark
Figure 9: Comparison of the OTTA performance on the ImageNet-C severity level $5$. The upper and bottom plots show the experiments conducted with batch size $16$ and $1$ based on the LayerNorm updating strategy, respectively. The dotted line represents the source-only performance.
When comparing performance across CIFAR-10-C, CIFAR-100-C, and ImageNet-C, it becomes apparent that ImageNet-C exhibits notably poorer performance overall. This could be attributed to the larger number of classes in the ImageNet-C dataset, as a similar trend is observed when comparing it with CIFAR-10-C and CIFAR-100-C.
[…]
Corruption Types. Varying domain gaps emerge across datasets when employing the same mechanism to generate corrupted data. Nevertheless, overall consistency remains, likely influenced by the training process of the source model. Additionally, corruptions representing natural domain gaps (e.g., brightness, zoom, snow) consistently appear simpler than noise-based corruptions (e.g., Gaussian noise, shot noise, etc.). Surprisingly, contrast exhibits a notably high error rate for both direct inference and most methods, indicating a substantial domain gap compared to other corruption types. In this case, ROID appears to excel in addressing it. In contrast, brightness appears to have the smallest domain gap with the source data.
Adaptation Strategy. As depicted in Fig. 9, when the batch size is set to 16, SAR, Conj-CE, RoTTA, and ROID outperform the source-only model in terms of mean error. In contrast, Tent, MEMO-LN, and CoTTA-LN demonstrate significantly poor results. The error rate within each domain exhibits a similar trend. Notably, Conj-CE, which conducts an additional inference of each batch for final prediction compared with Tent, markedly surpasses Tent across most domains and in mean error. This suggests a significant inter-batch shift in ImageNet-C, such as class differences, indicating that strategies based on optimizing the current batch for predicting the next batch (such as Tent) may be less effective.
4.2.4 On CIFAR-10-Warehouse Benchmark
In our study, the CIFAR-10-Warehouse dataset emerges as an indispensable resource for assessing OTTA methods, perfectly aligning with the CIFAR-10 label set to facilitate comprehensive comparisons across a spectrum of distribution shifts. Our detailed examination of two distinct domains within this dataset — real-world shifts and diffusion synthesis shifts — aims to probe deeper into the resilience and adaptability of OTTA methods amidst diverse distribution challenges.
Google split. Diverging from previous CIFAR-10 variances, which primarily integrated artificially induced corruptions (i.e., CIFAR-10-C) or relied on sample differences (i.e., CIFAR 10.1), the Google split of the CIFAR-10-Warehouse offers an unparalleled perspective. Sourced from Google search queries, this segment includes 12 subdomains that exhibit a range of color variations within the CIFAR-10 labels. It constructs an essential benchmark to measure contemporary OTTA methods’ proficiency against real-world distribution shifts.
Figure 10: Comparison of the OTTA performance on the Google split of CIFAR-10-Warehouse. The upper and bottom plots show the experiments conducted with batch size $16$ and $1$ based on the LayerNorm updating strategy. The dotted line represents the source-only performance.
Cross-dataset Comparison: Real-world v.s. Corruptions. Comparing the empirical results for both the Google split and CIFAR-10-C, a notable difference is that the real-world shift presented in CIFAR-10-Warehouse exhibits a smaller domain gap globally, as indicated by the performances of direct inference. Concerning subdomains, the real-world shifts also demonstrate lower variance compared to corruption domains. We can infer that dealing with attacks or noises in a streaming manner is the most challenging adaptation task, as these two datasets are based on the same source pre-trained model.
Batch Sizes. Regarding the differences in batch sizes depicted in Fig. 10, it is observed that when the batch size is 16, five OTTA methods based on LayerNorm updating either match or surpass the performance of direct inference. This observation suggests that the LayerNorm updating strategy is generally effective in handling real-world shifts. However, when the batch size is reduced to 1, methods that solely rely on entropy minimization, such as Tent and Conj-CE, experience performance degradation across most domains. This decline may be attributed to the instability of optimization for single-sample batches. Additionally, the fact that ROID performs similarly to corruption-based datasets indicates its sensitivity to changes in batch size.
Adaptation Strategy. RoTTA, TAST and SAR demonstrate exceptional stability regardless of batch sizes, which is also shown in CIFAR-100-C. Furthermore, ROID exhibits a smaller margin for surpassing performance compared to corruption-based datasets, indicating its robustness against data noise.
[…]